Grade 12th passMechanics

Speed of two identical cars are u and 4u at the specific instant.The ratio of the respective distance in which the two cars are stopped from that instant is1) 1:12)1:43)1:84)1:16

Dibyasingha Behera
9 Years agoGrade 12th pass

1 Answer

Sheetal kumari
9 Years ago
As you know. Stopping distance of a car is proportional to the square of velocity and stopping distance of a car depends only on it velocity and acceleration s=u^2/2a thus the ratio of respective distance in which it gets stopped is the square of the ratio of their velocity which is S1/S2=u^2/(4u)^2=u^2/16u^2=1/16 and thus your fourth option is correct Please kindly approved me if you clearly understand